# A Unified Framework for Input Feature Attribution Analysis

### v1

@@ 2.4 Simulatability
Simulatability was initially proposed to measure how accurately humans can predict a model’s outputs based on its explanations Chen et al. (2023). Previous studies have demonstrated that Simulatability can be approximated using an automated agent model as a surrogate for human understanding Pruthi et al. (2022). Given the established positive correlation between Simulatability and human evaluation of explanation utility from previous research, we integrate the Simulatability scores obtained from the agent model with different explanation types to approximate their utility for humans.
In prior research, an agent model $AM$, sharing the same architecture as the original model $M$, was introduced. $AM$ was trained to simulate $M$’s predictions $Y^{\prime}$ using post-hoc TokenEx explanations. During $AM$’s training phase, the top-$k$ explanations extracted for $M$ were provided as additional signals. In comparison, another agent model, $AM_{O}$, was trained without explanation guidance as a baseline on the same training set. During the testing phase, the simulation accuracy of both $AM$ and $AM_{O}$ over the shared dataset $D$ is calculated. The difference between their accuracies is interpreted as the explanation’s effect in enhancing the original model’s simulatability:
@@
To compare the simulation utility of different explanation types, we train a separate agent model, denoted as $AM_{E}$, for each explanation type $E$, and calculate the corresponding simulation performance on the common test set. For a fair comparison across the different explanation types, we first ensure top-$k_{E}$ explanations are presented for assisting the agent’s training for explanations type $E$, following the same method stated in Section §2.2. This approach ensures that each model is exposed to the same quantity of explanation information, specifically the number of unique tokens from the different explanation types.
During the training phase of the agent model, we introduce the explanations of type $E$ respectively into the learning of $AM_{E}$; we supplement the input sequence $x$ with top-$k_{E}$ explanations instead so that the agent model will be trained with the same mechanism whether the explanations are provided or not, and each training instance will contain the input sequence $x_{E}$ and golden label $Y^{\prime}$ which is predicted by the original model $M$. Specifically, we examine two different ways of presenting the explanations as part of the original input sequence, named $I_{Symbol}$ and $I_{Text}$; all aim to ensure the explanations of different types are inserted similarly. See the details of insertion forms in Appendix A.
At test time, the F1 scores of all agent models – $AM_{E}$ and $AM_{O}$, over the shared dataset $D$ are calculated. Note that the explanations of type $E$ are also provided to the input for the unseen instances for agent model $AM_{E}$. The difference between each agent model $AM_{E}$ and $AM_{O}$ on simulation performance is the final metric indicating how much this specific type of input feature explanation enhances the model’s simulatability, noted as $RSF_{E}$.
@@ 2.5 Complexity
Feature attribution explanations are designed to aid human understanding of a model’s reasoning over specific instances. However, since humans have a limited capacity to process large amounts of information simultaneously, these explanations need to be easy to comprehend. Even if we select only the top-$k$ attributions with the highest importance scores, they need to be distinctive as opposed to the attribution scores having a uniform distribution.
To measure this complexity of attribution scores, Bhatt et al. (2021) employ entropy Rényi (1961) over the attribution scores of all the produced highlight explanations. Higher entropy means different features have similar attribution scores, where the simplest explanation, with low entropy, would be concentrated on one feature. The calculation process is noted as follows, where m+n is the total number of generated TokenEx explanations, and all explanations are considered for the complexity score computation.
$P(x,p)=\frac{\left|a^{x}_{p}\right|}{\sum_{q=1}^{m+n}\left|a^{x}_{q}\right|}$ (17)
@@
Unified Complexity. Building on their approach, we also employ Entropy to measure the complexity of explanation attribution scores across different types of input feature explanations. To ensure a fair comparison, we maintain consistency in the size of the chosen explanation lists across all types for the same instance, denoted as $k_{x}$, as the number of generated TokenEx/TokenIntEx/SpanIntEx explanations vary for the same instance. The complexity score of the top-$k_{x}$ explanation list under type $E$, denoted as $CL_{E}(x,k_{x})$, is calculated as:
$P_{E}(x,k_{x},i)=\frac{\left|a^{x}_{i}\right|}{\sum_{j=1}^{k_{x}}\left|a^{x}_{j}\right|}$ (19)
$CL_{E}(x,k_{x})=-\sum_{i=1}^{k_{x}}P_{E}(x,k_{x},i)\ln(P_{E}(x,k_{x},i))$ (20)
where $a^{x}_{i}$/$a^{x}_{j}$ represent the $i$/$j$th highest attribution score from the explanation set for $x$. $P_{E}(x,k_{x},i)$ is the fraction of the attribution score of the $i^{th}$ most important explanation compared to the sum of the scores within all the top $k_{x}$ chosen explanations.
Finally, by applying this calculation of the complexity score regarding explanation attribution lists for type $E$ to all instances, we obtain the complexity score of each explanation type as $CL_{E}$.
$CL_{E}=\frac{\sum_{x\in D}^{\left|D\right|}CL_{E}(x,k_{x})}{\left|D\right|}$ (21)
Notably, we calculate $k_{x}$ using the number of SpanIntEx explanations within the instance, as the number of explanations of this type is unknown until they are generated, as discussed in §2.1.
@@ 3.1 Datasets
We select the natural language inference dataset SNLI (Bowman et al., 2015), where instances consist of a premise, a hypothesis, and a label $y\in\{entailment,neutral,contradiction\}$. Additionally, we select the fact checking dataset FEVER (Thorne et al., 2018), where instances consist of a claim, an evidence, and a label $y\in\{entailment,neutral,contradiction\}$. For generating the input feature explanations, we sample 4000 instances from each train, dev, and test set due to the computational requirements especially pronounced for Shapley-based explanations Atanasova et al. (2020). To evaluate agreement with human annotations, we describe in Appendix B the datasets used, which contain various types of input feature explanations for SNLI and FEVER annotated by humans.
@@ 3.2 Input Feature Explanation Methods
To generate importance scores, as techniques for TokenEx we select three common explanation techniques – Shapley (Lundberg and Lee, 2017), Attention (DeYoung et al., 2020), and Integrated Gradients (IG, Sundararajan et al. (2017)). For TokenIntEx, we employ Bivariate Shapley Masoomi et al. (2023), Attention (Clark et al., 2019), and Layer-wise Attention Attribution (Ye et al., 2021). Following Ray Choudhury et al. (2023), we apply the Louvain algorithm (Blondel et al., 2008) on top of each of the three selected TokenIntEx to generate importance scores for SpanIntEx, where the importance score of each span interaction is averaged over the importance scores of the token interactions within it.
Notably, the techniques applied for generating TokenIntEx are the bivariate version of the techniques used for generating TokenEx; e.g. Layer-wise Attention Attribution uses IG. Furthermore, SpanIntEx are generated from importance scores of TokenIntEx. We will refer to Shapley, Attention, and IG as the explanation base types used for generating all types of input feature explanations for brevity. See the details of generating all input feature explanations with different explainability techniques in Appendix D.
@@ 3.3 Models
We use the BERT-base-uncased model (Devlin et al., 2019) with 12 encoder layers, and the BART-base-uncased model (Lewis et al., 2020) with 6 encoder and 6 decoder layers, as common representatives of the encoder and the encoder-decoder Transformer architecture Vaswani et al. (2017). Our model choice is particularly influenced by the substantial computational requirements of the input feature explanations, especially pronounced for Shapley Atanasova et al. (2020). Additionally, our choice is guided by the need to directly access the models’ internals for generating IG and Attention-based explanations. Furthermore, while our framework currently utilizes the said models, it is designed to be easily adaptable to other models or newly developed explainability techniques, provided that there are more robust computational resources available.
We fine-tune the base models on SNLI and FEVER and use them to generate explanations. Their performance on the corresponding test sets is shown in Appendix C. For the Simulation property, we follow existing work Fernandes et al. (2022); Pruthi et al. (2022) and train simulator agent models (see §2.4) with the same architectures as the base ones. Following Fernandes et al. (2022), we split the test set into train/dev/test for the training of the agent model that simulates the original model’s predictions.
@@ 4 Results and Discussion
We now present the results of our unified evaluation framework (§2) illustrated in Figure 2. Our results include span the input feature explanations SpanIntEx, TokenIntEx, and TokenEx  (§3.2), two models (§3.3), two datasets (§3.1), and three base explanation techniques per input feature explanation type (§3.2). For Simulatability, we select the results of $I_{sym}$, as this form avoids repeating the input text and increasing the input size substantially. For Agreement with Human Annotation, we select the results of Token-level agreement as they are present for all types of explanations. Appendix E lists all detailed results per property.
@@ 4.1 Faithfulness
Unified Comprehensiveness. Across both datasets and models, TokenEx and TokenIntEx are identified as the most comprehensive explanation types, achieving the highest scores in 7/12 and 5/12 cases, respectively. SpanIntEx, designed to enhance the semantic coherence of interactive explanations by including additional context, often incorporates tokens that do not directly contribute to the model’s prediction, thus explaining its lower comprehensiveness scores. Compared to the random baseline, TokenEx and SpanIntEx always outperform it, while TokenIntEx mostly underperforms it when based on IG. Across the base explanation techniques, TokenEx performs best when based on Attention for BERT and on Shapley for BART, indicating that different base explanation techniques can perform better for different architectures. Both TokenIntEx and SpanIntEx show optimal performance when based on Shapley and Attention. Overall, the results indicate a stronger performance of Attention and Shapley over IG across all explanation types.
Unified Sufficiency. SpanIntEx ranks as the most sufficient explanation type in 7/12 cases, surpassing TokenEx, which performs well in only 3/12 cases. While contrary to SpanIntEx Comprehensiveness performance, we attribute this to the semantic coherence of the extracted top spans, which provide more meaningful information. Unlike TokenEx, which consistently outperforms the random baseline, TokenIntEx and SpanIntEx struggle to outperform it on FEVER, likely due to the longer input length compared to SNLI, posing more challenges for the explanations to accurately unveil the model’s internal processes. The results from different base explanation techniques show no clear trends, indicating a significant variability stemming from the specific dataset and model architecture.
@@ 4.2 Agreement with Human Annotation
In comparing token-level agreement, SpanIntEx and TokenIntEx show higher agreement scores with human annotations than TokenEx. Similarly, we find that SpanIntEx consistently achieves higher agreement with human interaction-level annotators, especially when based on Attention scores (see Appendix E). This indicates that SpanIntEx is more plausible to humans than the other types of explanations due to their enhanced semantics coherence. In contrast, TokenEx often scores lower than the random baseline. Moreover, considering SpanIntEx’s lower performance in terms of Comprehensiveness, there emerges a distinct trade-off between comprehensiveness and plausibility. Across the base explanation techniques, IG performs best for the FEVER dataset; IG and Attention perform best for SNLI. In addition, we find that for the interaction-level agreement, TokenIntEx and SpanIntEx perform worst when based on Shapley. The lower agreement results for Shapley compared to its better comparative performance on Comprehensiveness, once again indicate an existing trade-off between the two properties.
@@ 4.3 Simulatability
Our results show that SpanIntEx achieves the highest simulatability in 9/12 cases, helping the agent model accurately reproduce the original model’s predictions. This again underscores the critical role of contextual information and enhanced semantic coherence provided by SpanIntEx explanations. Notably, providing SpanIntEx explanations to agents improves their ability to simulate the original model by up to 7.9 F1 points compared to scenarios without explanations. Among the base explanation techniques, IG consistently performs best for SpanIntEx; other techniques do not exhibit a clear trend.
@@ 4.4 Complexity
SpanIntEx and TokenEx generally achieve similar complexity scores, which consistently remain lower than those of TokenIntEx. This suggests that TokenEx and SpanIntEx generate more distinctive attribution scores, potentially making them easier for humans to understand. Regarding the base explanation techniques, Attention consistently yields the best complexity scores for BERT across all explanation types. There is however no clear trend for BART. Additionally, as detailed in Appendix E, TokenIntEx frequently underperforms compared to a random baseline, highlighting its complexity.
@@ 4.5 Overall
In sum, we find that while TokenEx and TokenIntEx generally provide more comprehensive insights, SpanIntEx performs better in Sufficiency due to its enhanced semantic coherence, and despite occasionally including irrelevant tokens (§4.1). We find that there exists a trade-off between comprehensiveness and the plausibility of explanations (§4.2), suggesting that the most faithful explanations, and in turn, the model’s actual inner workings can be less plausible to end users.
Furthermore, SpanIntEx significantly helps for Simulatability, allowing agents to accurately replicate model decisions (§4.3), which is crucial for practical applications. The Complexity analysis (§4.4) shows that SpanIntEx and TokenEx are easier to comprehend compared to TokenIntEx, underscoring the importance of selecting explanation techniques that cater to the specific needs of the dataset and model architecture.
@@ 5 Related Work
Input Feature Explanations. Considerable research exists on extracting explanations from input data. Methods like perturbation-based attribution (e.g., Shapley (Lundberg and Lee, 2017)), attention-based methods (e.g., Attention (Jain and Wallace, 2019; Serrano and Smith, 2019)), and gradient-based methods (e.g., Integrated Gradients (Sundararajan et al., 2017; Serrano and Smith, 2019)) are prevalent for highlighting individual tokens. As individual tokens lack interaction information and thus might be insufficient to explain the model, many attribution methods have been extended to bivariate forms (Masoomi et al., 2023; Janizek et al., 2021; Sundararajan et al., 2017; Ye et al., 2021) to capture input token interactions. More recent work has explored how groups of tokens collectively contribute to model reasoning (Ray Choudhury et al., 2023; Chen et al., 2021). Unlike other work where token groups might consist of tokens from arbitrary positions, Ray Choudhury et al. (2023) introduces a method to explicitly capture span interactions, enhancing the comprehensiveness of explanations by containing the entire spans.
Explanation Evaluation. For evaluating TokenEx explanations, DeYoung et al. (2020); Atanasova et al. (2020) propose metrics to measure how faithful explanations are to the model’s inner reasoning. They also propose to assess the plausibility of explanations to humans by measuring the agreement of TokenIntEx with human annotations. To assess the utility of explanations to humans, Pruthi et al. (2022) propose to use an agent model as a proxy for humans and evaluate whether explanations aid in model simulatability. Complexity Bhatt et al. (2021) measures the distribution of attribution scores of TokenEx and assesses whether the key tokens in token explanations are easily comprehensible to humans. To evaluate TokenIntEx  most works adopt the faithfulness or axiomatic/theoretical path  Tsang et al. (2020); Sundararajan et al. (2020); Janizek et al. (2021). Current work on evaluating SpanIntEx explanations has primarily focused on faithfulness Ray Choudhury et al. (2023). However, since SpanIntEx, TokenIntEx, and TokenEx explanations contain varying amounts of tokens, which e.g. affects the faithfulness test, this makes direct comparisons between different explanation types using existing metrics challenging. To our knowledge, no prior paper has involved all types of input feature explanations within a unified evaluation framework.

### v2

@@ 2.4 Simulatability
Simulatability was initially proposed to measure how accurately humans can predict a model’s outputs based on its explanations Chen et al. (2024); Hase et al. (2020). Previous studies demonstrated that Simulatability can be approximated using an automated agent model as a surrogate for human understanding Pruthi et al. (2022). Given the established positive correlation between Simulatability and human evaluation of explanation utility, we integrate the Simulatability scores obtained from an agent model with different explanation types to approximate their utility for humans.
Following existing work Hase et al. (2020), we train an agent model $AM$, sharing the same architecture as the original model $M$, to simulate $M$’s predictions $Y^{\prime}$ using produced explanations. During $AM$’s training phase, we extract the top-$k$ explanations and incorporate them in the input. In comparison, another agent model, $AM_{O}$, is trained without explanation guidance as a baseline on the same training set. During the testing phase, the simulation accuracy of $AM$ and $AM_{O}$ over the shared dataset $D$ is calculated. [^2] While existing work Hase et al. (2020) notes that incorporating natural language explanations in the testing phase could leak the predicted label, we use only input feature explanations that do not contain additional information. The difference between the accuracies is interpreted as the explanation’s effect in enhancing the simulatability of $M$:
@@
To compare the simulation utility of different explanation types, we train a separate agent model $AM_{E}$ for each explanation method $A_{E}$ and calculate the corresponding simulation performance on the common test set. For a fair comparison across the different explanation method types $A_{E}$, we first ensure top-$k_{E}$ explanations are presented for assisting the agent’s training for $A_{E}$, following Section 2.2. This ensures each model is exposed to the same quantity of unique tokens from different explanation types.
During the training of $AM$, we introduce the explanations from $A_{E}$ into the learning of $AM_{A_{E}}$; we supplement $x$ with top-$k_{A_{E}}$ explanations instead so that the agent model is trained with the same mechanism whether the explanations are provided or not, and each training instance will contain the input sequence $x_{A_{E}}$ and golden label $Y^{\prime}$ which is predicted by the original model $M$. Specifically, we examine two different ways of presenting the explanations as part of the original input sequence, $I_{Symbol}$ and $I_{Text}$ (see §B.); all aim to ensure the explanations of different types are inserted similarly.
At test time, the F1 scores of agent models $AM_{E}$ and $AM_{O}$ over $D$ are calculated:
@@
Note that the explanations from $A_{E}$ are also provided to the input for the unseen instances for agent model $AM_{E}$. The final simulation metric then indicates how much this specific type of input feature explanation enhances the model’s simulatability:
@@ 2.5 Complexity
Feature attribution explanations are designed to aid human understanding of a model’s reasoning over specific instances. However, since humans have a limited capacity to process large amounts of information simultaneously, these explanations need to be easy to comprehend. Even if we select only the top-$k$ attributions with the highest importance scores, they need to be distinctive as opposed to the attribution scores having a uniform distribution. Bhatt et al. (2021) propose to measure the complexity of a produced explanation with entropy Rényi (1961) over the attribution scores of all the produced explanations by $A_{\texttt{TokenEx}}$ method:
$P(x,p)=\left|a^{x}_{p}\right|/\sum_{q=1}^{m+n}\left|a^{x}_{q}\right|$ (16)
@@
$m+n$ is the total number of generated $A_{\texttt{TokenEx}}$ explanations, and all explanations are considered for the complexity score computation. Higher entropy means different features have similar attribution scores, where the simplest explanation, with low entropy, would be concentrated on one feature.
Unified Complexity. To ensure a fair comparison across different types of explanation methods $A_{E}$, we maintain consistency in the size of the chosen explanation lists across all $A_{E}$ for the same instance, denoted as $k_{x}$, as the number of generated $A_{\texttt{TokenEx}}$/$A_{\texttt{TokenIntEx}}$/$A_{\texttt{SpanIntEx}}$ explanations originally vary for the same $x$. The complexity score $CL_{A_{E}}(x,k_{x})$ of the top-$k_{x}$ explanation list under method $A_{E}$ is calculated as:
$P_{A_{E}}(x,k_{x},i)=\left|a^{x}_{i}\right|/\sum_{j=1}^{k_{x}}\left|a^{x}_{j}\right|$ (18)
$CL_{A_{E}}(x,k_{x})=-\sum_{i=1}^{k_{x}}P_{A_{E}}(x,k_{x},i)\ln(P_{A_{E}}(x,k_{x},i))$ (19)
where $a^{x}_{i}$/$a^{x}_{j}$ represent the $i$/$j$th highest attribution score from the explanation set for $x$.
The final complexity score is an average of $CL_{A_{E}}(x,k_{x})$ across all $x\in D$:
$CL_{E}={\sum_{x\in D}^{\left|D\right|}CL_{E}(x,k_{x})}/{\left|D\right|}$ (20)
Notably, $k_{x}$ is calculated from the number of explanations produced by method $A_{\texttt{SpanIntEx}}$ for $x$, which varies based on the span interaction extraction method and is known only after generation.
@@ 3.1 Datasets
We select the natural language inference dataset SNLI (Bowman et al., 2015), where instances consist of a premise, hypothesis, and a label $y\in\{entailment,neutral,contradiction\}$. Additionally, we select the fact checking dataset FEVER (Thorne et al., 2018; Atanasova et al., 2020b), where instances consist of a claim, evidence, and a label $y\in\{entailment,neutral,contradiction\}$.333 https://huggingface.co/datasets/copenlu/fever_gold_evidence We generate input feature explanations by sampling 4,000 instances from each train, dev, and test set, due to the high computational cost, particularly for Shapley-based explanations Atanasova et al. (2020a). For Agreement with Human Annotations property, we use existing human explanation annotations for SNLI and FEVER (see §C).
@@ 3.2 Input Feature Explanation Methods
To generate importance scores, we first select three common TokenEx techniques – Shapley (Lundberg and Lee, 2017), Attention (DeYoung et al., 2020), and Integrated Gradients (IG, Sundararajan et al. (2017)). For TokenIntEx, we employ Bivariate Shapley Masoomi et al. (2022), Attention (Clark et al., 2019), and Layer-wise Attention Attribution (Ye et al., 2021). Notably, the TokenIntEx techniques are the bivariate version of the techniques used for generating TokenEx; e.g. Layer-wise Attention Attribution uses IG.
Following Ray Choudhury et al. (2023), we apply the Louvain algorithm (Blondel et al., 2008) for each of the three selected TokenIntEx to generate importance scores for SpanIntEx methods, where the importance score of each span interaction is averaged over the importance scores of the token interactions within it. We will refer to Shapley, Attention, and IG as the explanation base types used for generating all types of input feature explanations for brevity. See §E for more details.
@@ 3.3 Models
We use the BERT-base-uncased model (Devlin et al., 2019) with 12 encoder layers, and the BART-base-uncased model (Lewis et al., 2020) with 6 encoder and 6 decoder layers, as common representatives of the encoder and the encoder-decoder Transformer architecture Vaswani et al. (2017). Our choice is particularly influenced by the substantial computational requirements of the input feature explanations, especially pronounced for Shapley Atanasova et al. (2020a). Additionally, our choice is guided by the need to directly access the models’ internals for generating IG and Attention-based explanations. Furthermore, while our framework currently utilizes the said models, it is designed to be easily adaptable to other models or newly developed explainability techniques, provided that there are more robust computational resources available.
We fine-tune the base models on SNLI and FEVER and use them to generate explanations. Their performance on the test sets is shown in §D. For the Simulation property, we follow existing work Fernandes et al. (2022); Pruthi et al. (2022) and train simulator agent models (§2.4) with the same architectures as the base ones. Following Fernandes et al. (2022), we split the test set into train/dev/test for the training of the agent model.
@@ 4 Results and Discussion
We now present the results of our unified evaluation framework (§2) illustrated in Fig. 2. They include explanation methods of types SpanIntEx, TokenIntEx, and TokenEx (§3.2), two models (§3.3), two datasets (§3.1), and three base explanation techniques per explanation type (§3.2). For Simulatability, we select the results of $I_{sym}$, as this form avoids repeating the input text and increasing the input size substantially. For Agreement with Human Annotations, we select the Token-level results as they are present for all explanation types. §F lists detailed results per property.
@@ 4.1 Faithfulness
Unified Comprehensiveness. Across both datasets and models, TokenEx and TokenIntEx are identified as the most comprehensive explanation types, achieving the highest scores in 7/12 and 5/12 cases, respectively. SpanIntEx, designed to enhance the semantic coherence of interactive explanations by including additional context, often incorporates tokens that do not directly contribute to the model’s prediction, thus explaining its lower comprehensiveness scores. Compared to the random baseline, TokenEx and SpanIntEx always outperform it, while TokenIntEx mostly underperform it when based on IG. Across the base explanation techniques, TokenEx performs best when based on Attention for BERT and on Shapley for BART, indicating that different base explanation techniques can perform better for different architectures. Both TokenIntEx and SpanIntEx show optimal performance when based on Shapley and Attention. Overall, the results indicate a stronger performance of Attention and Shapley over IG across all explanation types.
Unified Sufficiency. SpanIntEx ranks as the most sufficient explanation type in 7/12 cases, surpassing TokenEx, which performs well in only 3/12 cases. While contrary to SpanIntEx Comprehensiveness performance, we attribute this to the semantic coherence of the extracted top spans, which provide more meaningful information. Note that while Sufficiency is highly desirable, Comprehensiveness is not required in all downstream applications as end-users prefer simpler, more general explanations with fewer causes Thagard (1989). Unlike TokenEx, which consistently outperforms the random baseline, TokenIntEx and SpanIntEx struggle to outperform it on FEVER, likely due to the longer inputing challenges for the explanations to accurately unveil the model’s internal processes. The results from different base explanation techniques show no clear trends, indicating a significant variability stemming from the specific dataset and model architecture.
@@ 4.2 Agreement with Human Annotation
SpanIntEx and TokenIntEx show higher agreement scores with human annotations than TokenEx. Similarly, SpanIntEx consistently achieve higher agreement with human interaction-level annotators, especially when based on Attention scores (see §F). This indicates that SpanIntEx are more plausible to humans due to their enhanced semantic coherence. In contrast, TokenEx often score lower than the random baseline. Moreover, considering SpanIntEx’s lower performance in Comprehensiveness, there emerges a distinct trade-off between Comprehensiveness and Agreement with Human Annotations. Across base explanation techniques, IG performs best for FEVER; IG and Attention – for SNLI. In addition, we find that for the interaction-level agreement, TokenIntEx and SpanIntEx perform worst when based on Shapley. The lower agreement results for Shapley compared to its better results on Comprehensiveness again indicate an existing trade-off between the two properties.
@@ 4.3 Simulatability
Our results show that SpanIntEx achieve the highest simulatability in 9/12 cases, helping the agent model accurately reproduce the original model’s prediction. This again underscores the critical role of contextual information and enhanced semantic coherence provided by SpanIntEx. Notably, providing SpanIntEx to agents improves their ability to simulate the original model by up to 7.9 F1 points compared to without explanations. Among base explanation techniques, IG consistently performs best for SpanIntEx; other techniques do not exhibit a clear trend.
@@ 4.4 Complexity
SpanIntEx and TokenEx generally achieve similar complexity, which consistently remains lower than those of TokenIntEx. This suggests that TokenEx and SpanIntEx generate more distinctive attribution scores, potentially making them easier for humans to understand. Regarding the base explanation techniques, Attention consistently yields the best complexity scores for BERT across all explanation types. There is no clear trend for BART. Additionally, TokenIntEx frequently underperform the random baseline, highlighting its complexity (see §F).
@@ 4.5 Overall
In summary, we find that while TokenEx and TokenIntEx generally provide more comprehensive insights, SpanIntEx performs better in Sufficiency due to its enhanced semantic coherence (§4.1). This calls for better methods for generating SpanIntEx that are both comprehensive and sufficient. Additionally, there is a trade-off between Comprehensiveness and Agreement with Human Annotations (§4.2), suggesting that the most faithful explanations might be less plausible to end users. This highlights the need for advanced methods to boost both the Comprehensiveness and plausibility of SpanIntEx  possibly leveraging the advantage of TokenEx. Furthermore, SpanIntEx significantly improves simulatability by allowing agents to accurately replicate model decisions (§4.3), which is crucial in practice. Finally, the complexity analysis (§4.4) shows that SpanIntEx and TokenEx are potentially easier to comprehend than TokenIntEx when considering the importance score distribution.
Overall, our results highlight the differences between the different types of input feature explanations, with SpanIntEx outperforming others on most measures. As no one type performs best on all properties, we call for the development of combined methods that can leverage the strength of the different explanation types and potentially lead to an overall improvement of the explanation utility.
@@ 5 Related Work
Input Feature Explanations. Considerable research exists on extracting explanations for input data. Methods like perturbation-based attribution (e.g., Shapley (Lundberg and Lee, 2017)), attention-based methods (e.g., Attention (Jain and Wallace, 2019; Serrano and Smith, 2019)), and gradient-based methods (e.g., Integrated Gradients (Sundararajan et al., 2017; Serrano and Smith, 2019)) are prevalent for highlighting individual tokens (Atanasova et al., 2020a). As individual tokens might be insufficient to explain the model, many attribution methods have been extended to bivariate forms (Masoomi et al., 2022; Janizek et al., 2021; Sundararajan et al., 2017; Ye et al., 2021) to capture input token interactions. More recent work has explored how interactions between groups of tokens collectively contribute to model reasoning (Ray Choudhury et al., 2023; Chen et al., 2021). Unlike other work where token groups might consist of tokens from arbitrary positions, Ray Choudhury et al. (2023) explicitly capture span interactions, enhancing the comprehensiveness of explanations by containing the entire spans.
Automated Explanation Evaluation. For evaluating TokenEx, DeYoung et al. (2020); Atanasova et al. (2020a) propose metrics to measure how faithful explanations are to the model’s inner reasoning. They also assess the plausibility of explanations to humans by measuring the agreement of TokenIntEx with human annotations. To assess the utility of explanations to humans, Pruthi et al. (2022) propose to use an agent model as a proxy for humans and evaluate whether explanations aid in model simulatability. Complexity Bhatt et al. (2021) measures the distribution of attribution scores of TokenEx and assesses whether the key tokens in token explanations are easily comprehensible to humans. To evaluate TokenIntEx  most works adopt the faithfulness or axiomatic/theoretical path  Tsang et al. (2020); Sundararajan et al. (2020); Janizek et al. (2021). Current work on evaluating SpanIntEx has primarily focused on faithfulness Ray Choudhury et al. (2023). However, since SpanIntEx, TokenIntEx, and TokenEx contain varying amounts of tokens, which, e.g., affects the faithfulness test, this makes direct comparisons between different explanation types using existing metrics challenging. To our knowledge, no prior paper has involved all types of input feature explanations within a unified evaluation framework.